OpenStudy (anonymous):

calculate the derivative of y with respect to x. (x^3)y+3x(y^3)= x + y

OpenStudy (anonymous):

Do you know about the product rule of derivative??

OpenStudy (anonymous):

Yes i do and applied it, but for some reason the answer i've been getting is not correct.

OpenStudy (anonymous):

Can you show me the entire steps so that I can fine the mistake where you are doing it??

OpenStudy (anonymous):

*find..

OpenStudy (anonymous):

\[\frac{ dy }{ dx }[3x^2y+\frac{ dy }{ dx }x^3]+[xy^3+3y^2\frac{ dy }{ dx }3x^2] = 1 +\frac{ dy }{ dx }\]

OpenStudy (anonymous):

that's how i did it, i'm not sure if i made an algebra mistake

OpenStudy (anonymous):

Why are you using dy/dx with the first term??

OpenStudy (anonymous):

what are you referring to? I use dy dx because it said to find the derivative of y in respect to x

OpenStudy (anonymous):

Let us take the first term first.. Do it slowly.. So we have : \(x^3y\).. Find its derivative..

OpenStudy (anonymous):

\[x^3 y = x^3 \cdot \frac{dy}{dx} + 3x^2 \cdot y\] Isn't it??

OpenStudy (anonymous):

well it's basically what i wrote early except i switched the order of the number i took the derivative of.

OpenStudy (anonymous):

earlier*

OpenStudy (anonymous):

But you have attached dy/dx with that term.. Okay I understand your feelings that you are just showing that is the derivative only..

OpenStudy (anonymous):

But in actual attaching dy/dx with that term will make your question wrong..

OpenStudy (anonymous):

Well since i'm not sure how to solve it, would you be able to lay out the steps in which it is correctly solved?

OpenStudy (anonymous):

And derivative of second term will be: \[3xy^3 = 3[x \cdot 3y^2 + y^3]\]

OpenStudy (anonymous):

Just wait... To arrange steps it will take time.. Just few minutes..

OpenStudy (anonymous):

but what about the y prime? am i not suppose to include it?

OpenStudy (anonymous):

yes I forgot there, y' will be there with y^2..

OpenStudy (anonymous):

\[\frac{d}{dx}(x^3y) + \frac{d}{dx}(3xy^3) = \frac{d}{dx}(x + y)\] So this we have to solve as our question is saying..

OpenStudy (anonymous):

See, firstly you know fully about the product rule and how to apply it, then it will be very simple..

OpenStudy (anonymous):

Internet is not working properly..

OpenStudy (anonymous):

alright so what we'll have after the product rule is going to be \[[3x^2y+y'x^3]+3[y^3+3y^2y'x]=1+y'\]

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

So let us go further.. \[\implies x^3y' + 3yx^2 + 3(x \cdot 3y^2 \cdot y' + y^3) = 1 + y'\]

OpenStudy (anonymous):

Yeah that is right..

OpenStudy (anonymous):

and then i subract the 3yx^2 to the right and then subtract y' from the left right?

OpenStudy (anonymous):

and then i'm stuck after that

OpenStudy (anonymous):

Sorry, my internet connection is somewhat slow..

OpenStudy (anonymous):

Now proceeding further we can separate y' terms and normal terms like this : \[3x^2y + 3y^3 - 1 = y' + x^3y' + 9xy^2y'\]

OpenStudy (anonymous):

Getting this step?? I am just separating normal terms on left hand side and y' terms on right..

OpenStudy (anonymous):

I am expecting a reply from your side @duh_itsme

OpenStudy (anonymous):

oh yes i'm sorry i do

OpenStudy (anonymous):

and the i factor out the y' right?

OpenStudy (anonymous):

Yep, so go ahead.. This step you will do..

OpenStudy (anonymous):

wait, for the multiplication rule of the second part, shouldn't it be 27xy^2y'? since you have to distribute in the 3, 3x, and 3y^2?

OpenStudy (anonymous):

oh never mind, just ignore what i just said.

OpenStudy (anonymous):

Ha ha ha.. Are you okay?? Can you go further??

OpenStudy (anonymous):

Just factor out y' from RHS and show me where we have arrived..

OpenStudy (anonymous):

oh yeah, hahahah i figured it out. My brain just went wacko for a moment. lol!

OpenStudy (anonymous):

y'(-1+x^3+9xy^2) is what i got after factoring out y'

OpenStudy (anonymous):

I just made a mistake and I think you are right..

OpenStudy (anonymous):

now...i'm confused......

OpenStudy (anonymous):

Are you doing it your way or according to my equation??

OpenStudy (anonymous):

your way

OpenStudy (anonymous):

Okay then there are two minus mistake..

OpenStudy (anonymous):

\[3x^2y + 3y^3 - 1 = y' -x^3y' - 9xy^2y'\]

OpenStudy (anonymous):

This should be the equation .. Have you got how??

OpenStudy (anonymous):

Bringing positive terms from LHS will make them Negative on RHS.. Getting??

OpenStudy (anonymous):

I actually got something different.

OpenStudy (anonymous):

Can you show me??

OpenStudy (anonymous):

give me a moment please.

OpenStudy (anonymous):

Take your time.. And I am really very sorry, I am just making mistakes while solving, that is not good for me as well as for the person like you whom I am teaching..

OpenStudy (anonymous):

it's alright, you're still helping me. :P that's what counts

OpenStudy (anonymous):

Now let us just concentrate on it and solve it out.. Can you show me what you get and take your time, I am in no hurry..

OpenStudy (anonymous):

so what i got was y'x^3+9xy^2y'-y' = -3x^2y-3y^3+1

OpenStudy (anonymous):

i guess it's the same as yours except the negative is distributed differently.

OpenStudy (anonymous):

I was thinking the same. the terms I gathered on LHS, you are just gathering them on RHS, and it is no wrong.. So you are right dear..

OpenStudy (anonymous):

Yep.. Now factor out y' from this..

OpenStudy (anonymous):

yeah and then divide the equation on both sides right?

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

\[y'(x^3 + 9xy^2 - 1) = 1 - 3x^2y - 3y^3\] Now just divide and find \(y'\)..

OpenStudy (anonymous):

You are welcome dear..